# Controllable non-Hermitian qubit-qubit Coupling in Superconducting quantum Circuit

### v1

@@ Abstract
With a high-loss resonator supplying the non-Hermiticity, we study the Energy level degeneracy and quantum state evolution in tunable coupling superconducting quantum circuit. The qubit’s effective energy level and damping rate can be continually tuned in superconducting circuit, and the positions and numbers of level degenerate points are controllable. The efficient of quantum state exchange and the asymmetry of quantum state evolution can be tuned with non-hermitian and nonreciprocal coupling between two qubits. The controllable non-Hermiticity provides new insights and methods for exploring the unconventional quantum effects in superconducting quantum circuit.
@@ I Introduction
The non-Hermitian system possessing complex energy spectrum, the real and imaginary parts respectively label the energy levels and loss of system, while the degenerate point of energy levels are known as the Exceptional points[3, 4, 5, 6, 7]. The system close to exceptional points contain rich unconventional physical effects, including the ultralow driving threshold chaos[23], high sensitivity metrology[9, 10], high-dimensions skin effect[11, 12], Chiral Heat Transport[13] and so on. The exceptional point have been studied in single ion system[14, 15], optical microcavity[16, 17, 18, 19, 20, 21, 22, 24], and optomechanical system[25, 26, 27, 28], magnonic system[29, 30, 31], Hybrid Quantum System[23], single superconducting qubit[32, 33], and so on. Some theoretical work have studied the PT-symmetry in superconducting circuit with the single qubit coupling to the resonatorcite[34, 35]. However, it is still challenging to experimentally observe such non-Hermitian in multi-qubits superconducting circuit due to the difficulties in realizing controlled gain and loss.
@@
By preparing one qubit in the excited state and another in the ground states, the free evolution of quantum states under non-hermite coupling can be tuned with the J-C coupling phases. Even setting the identical paramters, the evolutions of two qubits’ quantum states are different from different coupling directions. The non-reciprocal couplings between in superconducting circuit can be used to develop the integrated unidirectional transmission devices including the microwave circulator and isolator.
@@ II Effective Non-Hermitian Circuit
As shown in Fig. 1, we study a superconducting circuit consisting of two Xmon qubits coupling to a common transmon coupler and a high-loss superconducting resonator. The transmon coupler and the high-loss resonator can induce the hermite and non-hermite indirect interaction between two qubits. Defining $\omega_{k}$ as the transition frequencies of superconducting artificial atoms, and $\sigma^{(k)}_{z}$ and $\sigma^{(k)}_{\pm}$ are the respective pauli-Z and ladder operators, where $k=1,2,c$ label the qubit 1, qubit 2 and transmon coupler, respectively.
@@
The free Hamiltonian for the superconducting circuit can be written as $H_{0}=\omega_{a}a^{\dagger}a+\sum_{k=1,2,c}(\omega_{k}/2)\sigma^{(k)}_{z}$. $\omega_{a}$ is the resonant frequency of resonator, and the corresponding transition frequency of superconducting artificial atoms are defined as $\omega_{k}$, where $k=1,2,c$ label the qubit 1, qubit 2, and the transmon coupler, respectively. $a^{\dagger}$ and $a$ are the creation and annihilation operators for the resonator photons, and $\sigma^{(k)}_{z}$ is the Pauli-Z operator of superconducting artificial atom. The Hamiltonian for the direct qubit-qubit interaction is $H_{xy}=g_{xy}[\sigma^{(1)}_{+}\sigma^{(2)}_{-}+\sigma^{(2)}_{+}\sigma^{(1)}_{-}]$, where $g_{xy}$ label the direct coupling strength between two qubits. We defined the $\sigma^{(k)}_{\pm}$ as the ladder operators of artificial atoms. The Hamiltonian for the interaction between coupler and qubit-$j$ can be written as $H^{j}_{qc}=g_{j}[\sigma^{(c)}_{+}\sigma^{(j)}_{-}+\sigma^{(j)}_{+}\sigma^{(c)}_{-}]$, where $g_{j}$ is the coupling strength between qubit-$j$ and transmon coupler, with $j=1,2$.
@@
Figure 1: (Color online) The superconducting circuit consisting of two Xmon qubits coupling to a common transmon coupler and a high-loss resonator. The transition frequency of superconducting artificial atom is $\omega_{k}$, where $k=1,2,c$ label the qubit 1, qubit 2, and transmon coupler, the resonant frequency of resonator is $\omega_{a}$. The capacitance type interaction is dominated between arbitrary two devices.
@@
In this article, the frequency detuning microwave is applied on qubit to create amplitude and phase tunable J-C (Jaynes Cummings) interaction between qubit and resonator[36, 37, 38]. The qubits and resonator are in the dispersive coupling regimes, that is $\lambda^{(0)}_{j}\ll|\Delta_{ja}|,|\Delta_{ja}+\alpha_{j}|$, where $\lambda^{(0)}_{j}$ labels the bare J-C coupling amplitude between first-excited state of superconducting artificial atom-$j$ and the fundamental mode of resonator and their frequency detuning is $\Delta_{ja}=\omega_{j}-\omega_{a}$, while $\alpha_{j}$ describe the anharmoncity of qubit-$j$. After injecting a single tone microwave pulse with amplitude $\Omega_{j}(t)$ and phase $\theta_{j}(t)$, thus J-C interaction between resonator and qubit-$j$ can be obtained as $\lambda_{j}\exp{(\theta_{j})}$, with $\lambda_{j}=(1/\sqrt{2})\lambda^{(0)}_{j}\alpha_{j}|\Omega_{j}|/[\Delta_{ja}(\Delta_{ja}+\alpha_{j})]$[36, 37]. The amplitude $\lambda_{j}$ and phase $\theta_{j}$ can be tuned by the input microwave pulse, thus the Hamiltonian for amplitude and phase tunable J-C coupling can be written as $H^{j}_{qr}=[\lambda_{j}\exp{(-i\theta_{j})}a^{\dagger}\sigma^{(j)}_{-}+\lambda_{j}\exp{(i\theta_{j})}\sigma^{(j)}_{+}a]$. The interaction between resonator and transmon coupler is very weak and can be neglected, thus the Hamiltonian for the superconducting circuit in Fig. 1 can be obtained as $H=H_{0}+H_{xy}+\sum_{j}H^{j}_{qc}+\sum_{j}H^{j}_{qr}$.
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 2: (Color online) Controllable levels degeneracy in non-hermitian superconducting circuit. (a) Effective energy levels and (b) damping rates as the function of effective coupling $g_{e}$: $\Delta\theta=0.500\pi$ (blue curves), $0.501\pi$ (red curves), and $0.502\pi$ (black curves). The two qubits are assumed in the ground states $\langle\sigma^{j}_{z}\rangle=-1$, with $j=1,2$. (c) Effective energy levels as the function of J-C coupling amplitude $\lambda_{j}$: $\Delta\theta=0.500\pi$, $\Gamma_{2}=\Gamma_{1}$, $g_{e}/2\pi=1.137$ MHz, and $\gamma_{a}/2\pi=65$ MHz (blue curves); $g_{e}/2\pi=1.224$ MHz (green curves); $\gamma_{a}/2\pi=66$ MHz (red curves); $\Gamma_{2}=0.99\Gamma_{1}$(black curves). $\langle\sigma^{(1)}_{z}\rangle=-1$, and other un-mentioned parameters on the green, red, and black curves are the same as that in blue curves. (d) Effective energy levels as function of qubit populations: $\Delta\theta=0.500\pi$, $\lambda_{j}/2\pi=11$ MHz, $g_{e}/2\pi=1.137$ MHz (blue curves); $\Delta\theta=0.5005\pi$, the other parameters same as blue curve (red curves); $g_{e}/2\pi=1.224$ MHz, $\lambda_{j}/2\pi=11$ MHz, other parameters same as blue curves (black curves). The population of two qubits are assumed oppositive, $\langle\sigma^{j}_{z}\rangle=(-1)^{(j+1)}\langle\sigma_{z}\rangle$) The other parameters in (a)-(d) are: $\omega_{a}/2\pi=4.475$ GHz, $\omega_{1}/2\pi=4.5$ GHz, $\omega_{2}/2\pi=4.505$ GHz, $\omega^{(max)}_{c}/2\pi=5.20$ GHz, $\gamma_{1}/2\pi=1.00$ MHz, $\gamma_{2}/2\pi=1.01$ MHz, $g_{xy}/2\pi=4.0$ MHz, $\lambda_{j}/2\pi=11$ MHz, $g_{1}/2\pi=30$ MHz, $g_{2}/2\pi=30.3$ MHz, and $\gamma_{a}/2\pi=65$ MHz.
In the rotating coordinates at resonator’s frequency $\omega_{a}$, the motion equations of resonator and qubits can be written as $\dot{a}=-\gamma_{a}a-i\sum_{j=1,2}\lambda_{j}\exp{(-i\theta_{j})}\sigma^{(j)}_{-}$ and $\dot{\sigma}^{(j)}_{-}=-\left[\gamma_{j}/2+i\left(\Delta_{ja}+g^{2}_{j}/\Delta_{jc}\right)\right]\sigma^{(j)}_{-}-i\lambda_{j}\exp{(i\theta_{j})}\sigma^{(j)}_{z}a-ig_{e}\sigma^{(j)}_{z}\sigma^{(j+1)}_{-}$, with $\Delta_{ja}=\omega_{j}-\omega_{a}$ and $j+1=1$ for $j=2$. In the case of $\gamma_{a}\gg g_{j},\lambda_{j}\gg\gamma_{j}$ and $|\Delta_{ja}|/\gamma_{a}\approx 1/3$, with the adiabatic approximation, we can get $a=-i(\lambda_{1}/\gamma_{a})\sum_{j=1,2}\lambda_{j}\exp{(-i\theta_{j})}\sigma^{(j)}_{-}$. Substituting it into the motion equation of ladder operators, thus we get
@@
The Pauli-Z operator $\sigma^{(j)}_{z}$ describes the population of qubit-$j$ and its steady value can be tuned with the external microwave signals. Though coupling to a high-loss resonator, the decoherence time of qubit is still much longer than the durations for preparation and readout of single qubit’s quantum state (usually below 100 ns) especially for the tantalum-based superconducting qubit[44, 45].
Thus the population of qubit-$j$ can be considered as staying in the transient stability value $\langle\sigma^{(j)}_{z}\rangle$ during the preparations and measurements of qubits’ quantum states, and its value can be tuned by the qubit-coupler coupling strength, and also the J-C coupling phases and amplitudes. Thus the effective Hamiltonian for non-Hermitian qubit-qubit coupling can be written as
@@
where $\Delta\theta=\theta_{1}-\theta_{2}$, $\Delta^{\prime}_{ja}=\Delta_{ja}+g^{2}_{j}/\Delta_{jc}$, and $\Gamma_{j}=\gamma_{j}-\lambda^{2}_{j}\langle\sigma^{(j)}_{z}\rangle/\gamma_{a}$, with $j=1,2$. The non-diagonal terms in Eq.(3) label total effective qubit-qubit coupling strength in the nonhermite superconducting circuit. The two Eigen-modes for the non-Hermitian Hamiltonian $H^{\prime\prime}$ can be calculated as
@@
The real part $\Re{(\omega_{\pm})}$ corresponds to the effective energy levels of two qubits, while the imaginary part $\Im{(\omega_{\pm})}$ describes effective damping rates. The values of $R$ and $I$ affect the separations of complex energy spectrum between two Eigen-modes, and the signs of $R$ decide the level repulsions or attractions, while the sign of $I$ describes the relative size of two effective damping rates.
@@ III Controllable Non-Hermitian coupling
The experimental tunable parameters in superconducting circuit of Fig. 1 include the J-C coupling phases ($\theta_{j}$) and amplitudes $\lambda_{j}$, effective qubit-qubit coupling $g_{e}$, and population $\sigma^{(j)}_{z}$ of qubit $j$, which can be continually tuned to find the available parameters for exceptional point. The $\sqrt{R+iI}$ becomes a pure imaginary number($\Delta E_{q}=0$) in the case of $R<0$ and $I=0$, so the energy levels of two qubits can merge as one and the exceptional point should appear. The two qubits will possess misaligned energy levels even for a small nonzero value of $I$, and the exceptional point should be absent[28].
@@
By setting $\langle\sigma^{(j)}_{z}\rangle=-1$ ($\Gamma_{1}-\Gamma_{2}\neq 0$) and J-C coupling amplitude $\lambda_{j}/(2\pi)=11$ MHz, the effects of J-C coupling phases on the effective energy levels and damping rates of two Eigen-modes are shown in Figs. 2(a) and  2(b), respectively. The separations of two qubits’ effective energy levels (or damping rates) becomes narrower (or wider) in certain parameter regimes corresponding to the the energy level attraction. To see clearer, the curves in blue and black color shift up -0.5 MHz or 0.5 MHz in Fig. 2(a), respectively. As shown in the inset figure of Fig. 2(a), there is a degenerate points on the pairs of two red or black color curves for $\Delta\theta\neq\pi/2$. In the case of $\Delta\theta=\pi/2$, the first term in the right side of Eq.(6) is always zero, since $\Delta^{\prime}_{1a}-\Delta^{\prime}_{2a}\neq 0$ in parameter area of drawing, so there is no level degenerate point on the blue color curves in Fig. 2(a). The scattering curves for damping rates of two Eigen-mode are shown in Fig. 2(b), to see clearer, the curves in red and black color shift up 0.1 MHz and 0.2 MHz, respectively. The two branch of effective damping rates exchange with each other in two blue or black color curves in Fig. 2(b), and the corresponding exchange positions coincides well with the corresponding locations of level degenerate points in Fig. 2(a), which should indicate the positions of Exceptional points.
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Figure 3: (Color online) Controllable quantum state evolutions in energy level attraction regimes. (a) Differences in effective energy level ($\Delta E_{q}$) and (b) in effective damping rate ($\Delta\gamma_{q}$) as the functions of phase difference $\Delta\theta$ and coupling strength $g_{e}$. The two qubits are assumed as in the ground states, $\langle\sigma^{(j)}_{z}\rangle=-1$ with $j=1,2$. The free evolutions for excited state occupation populations for qubits (c) starting in excited state or (d) ground state for the phase differences: $\Delta\theta=0$ (blue-solid curves); $\Delta\theta=\pi/3$ (green-dashed curves); $\Delta\theta=5\pi/12$ (red-dotted curves); $\Delta\theta=\pi/2$ (black-dash-dotted curves). A small effective coupling ($g_{e}/2\pi=-80$) kHz to realize the energy level attractions, and the other parameters in (a)-(d) are: $\omega_{a}/2\pi=4.475$ GHz, $\omega_{1}/2\pi=4.5$ GHz, $\omega_{2}/2\pi=4.505$ GHz, $\omega^{(max)}_{c}/2\pi=5.20$ GHz, $\gamma_{1}/2\pi=1.00$ MHz, $\gamma_{2}/2\pi=1.01$ MHz, $g_{xy}/2\pi=4.0$ MHz, $\lambda_{j}/2\pi=11$ MHz, $g_{1}/2\pi=30$ MHz, $g_{2}/2\pi=30.3$ MHz, and $\gamma_{a}/2\pi=65$ MHz.
@@
The effects of J-C coupling amplitudes on energy levels of two qubit are studied in Fig. 2(c). In the case of $\Delta\theta=\pi/2$ and $\Gamma_{1}-\Gamma_{2}=0$, the first and second terms in Eq.(6) are always zeroes, so the energy level degeneracy appear for arbitrary value of $\lambda_{j}$ (energy level attraction regimes) in the blue curves. The change for the values of $g_{e}$ can shift the position of the Exception point and also the value of energy levels as indicated by the green curves (compared with the blue curves). If the $\Gamma_{1}-\Gamma_{2}\neq 0$, because of no solution for $\Delta^{\prime}_{1a}-\Delta^{\prime}_{2a}\neq 0$ in the parameter area of drawing, so the energy level degenerate points does not appear in the red and black curves. As indicated by the motion equation of $\sigma^{(j)}_{z}$, the effective energy levels of two qubits can be tuned by changing the quantum states of qubits.
@@
The free evolutions of quantum states under non-hermitian coupling can be calculated with the semiclassical limit method by disregarding the quantum jump terms[33, 49, 50]. By preparing the qubit 1 in the excited states and qubit 2 in the ground states, the time evolution of occupation probabilities for qubits’ excited state can be calculated with the semiclassical master equation $\dot{\rho}=-i(H^{\prime\prime}\rho-\rho H^{\prime\prime})$, where the $\rho$ is the density operator. To keep the superconducting circuit in the level attraction regimes, the effective qubit-qubit coupling $g_{e}$ is tuned to a small value, $g_{e}/2\pi=-80$ kHz. The time evolution of occupation probabilities for qubit-$j$ starting in excited ($P^{(se)}_{j}$) or ground ($P^{(sg)}_{j}$) state is shown in Fig. 3(c) or Fig. 3(d), respectively. The periodic oscillation of the curves originates from the nonzero frequency detuning $\Delta_{ja}$ in $H^{\prime\prime}$, and the effects of the J-C coupling phase can be reflected by the envelopes of the oscillating curves.

### v2

@@ Abstract
With a high-loss resonator supplying the non-Hermiticity, we study the energy level degeneracy and quantum state evolution in the tunable coupling superconducting quantum circuit. The qubits’ effective energy levels and damping rates can be continually tuned by the qubit-resonator coupling phase and amplitude, quantum states of qubits, and effective qubit-qubit coupling strength, thus the positions and numbers of level degenerate points are controllable. The efficient of quantum state exchange and the asymmetry of quantum state evolution can be tuned with non-hermitian and nonreciprocal coupling between two qubits. The controllable non-Hermiticity provides new insights and methods for exploring the unconventional quantum effects in superconducting quantum circuit.
@@ I Introduction
Non-Hermitian system possessing complex energy spectrum, the corresponding real and imaginary parts respectively label the energy levels and loss, and the degenerate point of energy levels are known as the exceptional points[1, 2, 3, 4, 5]. The system close to exceptional points contain rich unconventional physical effects, including the ultralow driving threshold chaos[29], high sensitivity metrology[7, 8], high-dimensions skin effect[9, 10], chiral heat transport[11] and so on. The exceptional point have been studied in single ion system[12, 13], optical microcavity[14, 15, 16, 17, 18, 19, 20, 21], and optomechanical system[22, 23, 24, 25], magnonic system[26, 27, 28], hybrid quantum system[29], single superconducting qubit[30, 31], and so on. Some theoretical work have studied the $\mathcal{P}\mathcal{T}$-symmetry in superconducting circuit with the single qubit coupling to the resonator[32, 33]. However, it is still challenging to experimentally observe such non-Hermitian in multi-qubits superconducting circuit due to the difficulties in realizing controlled gain and loss.
@@
By preparing one qubit in the excited state and another in the ground states, the free evolution of quantum states under non-Hermite interaction can be tuned with the resonator-qubit coupling phases. Even setting the identical parameters, the evolutions of two qubits’ quantum states are different from different coupling directions. The non-reciprocal couplings in superconducting circuit can be used to develop the integrated unidirectional transmission devices including the microwave circulator and isolator.
@@ II Effective Non-Hermitian Circuit
As shown in Fig. 1, we study a superconducting circuit consisting of two Xmon qubits coupling to a common transmon coupler and a high-loss superconducting resonator. The transmon coupler and the high-loss resonator can induce the hermite and non-hermite indirect interaction between two qubits. The free Hamiltonian for the superconducting circuit can be written as $H_{0}=\omega_{a}a^{\dagger}a+\sum_{k=1,2,c}(\omega_{k}/2)\sigma^{(k)}_{z}$. $\omega_{a}$ is the resonant frequency of resonator, and the corresponding transition frequency of superconducting artificial atoms are defined as $\omega_{k}$, where $k=1,2,c$ label the qubit 1, qubit 2, and the transmon coupler, respectively. $a^{\dagger}$ and $a$ are the creation and annihilation operators for the resonator photons, and $\sigma^{(k)}_{z}$ is the Pauli-Z operator of artificial atom. The Hamiltonian for the direct qubit-qubit interaction is $H_{xy}=g_{xy}[\sigma^{(1)}_{+}\sigma^{(2)}_{-}+\sigma^{(2)}_{+}\sigma^{(1)}_{-}]$, where $g_{xy}$ label the direct coupling strength between two qubits. We define the $\sigma^{(k)}_{\pm}$ as the ladder operators of artificial atoms. The Hamiltonian for the interaction between coupler and qubit-$j$ can be written as $H^{j}_{qc}=g_{j}[\sigma^{(c)}_{+}\sigma^{(j)}_{-}+\sigma^{(j)}_{+}\sigma^{(c)}_{-}]$, where $g_{j}$ is the coupling strength between qubit-$j$ and transmon coupler, with $j=1,2$.
@@
Figure 1: (Color online) The superconducting circuit consisting of two Xmon qubits coupling to a common transmon coupler and a high-loss resonator. The capacitance type interaction is dominated in the superconducting circuit.
@@
In this article, the frequency detuning microwave is applied on qubit to create amplitude and phase tunable J-C interaction between qubit and resonator[34, 35, 36]. The qubits and resonator are in the dispersive coupling regimes, that is $\lambda^{(0)}_{j}\ll|\Delta_{ja}|,|\Delta_{ja}+\alpha_{j}|$, where $\lambda^{(0)}_{j}$ labels the bare resonator-qubit coupling amplitude between first-excited state of superconducting artificial atom-$j$ and the fundamental mode of resonator and their frequency detuning is $\Delta_{ja}=\omega_{j}-\omega_{a}$, while $\alpha_{j}$ describe the anharmoncity of qubit-$j$. After injecting a single tone microwave pulse with amplitude $\Omega_{j}(t)$ and phase $\theta_{j}(t)$, thus resonator-qubit-$j$ interaction can be obtained as $\lambda_{j}\exp{(\theta_{j})}$, with $\lambda_{j}=(1/\sqrt{2})\lambda^{(0)}_{j}\alpha_{j}|\Omega_{j}|/[\Delta_{ja}(\Delta_{ja}+\alpha_{j})]$[34, 35]. The amplitude $\lambda_{j}$ and phase $\theta_{j}$ can be tuned by the input microwave pulse, thus the Hamiltonian for amplitude and phase tunable resonator-qubit coupling can be written as $H^{j}_{qr}=[\lambda_{j}\exp{(-i\theta_{j})}a^{\dagger}\sigma^{(j)}_{-}+\lambda_{j}\exp{(i\theta_{j})}\sigma^{(j)}_{+}a]$. The interaction between resonator and transmon coupler is very weak and can be neglected, thus the Hamiltonian for the superconducting circuit in Fig. 1 can be obtained as $H=H_{0}+H_{xy}+\sum_{j}H^{j}_{qc}+\sum_{j}H^{j}_{qr}$.
@@
Figure 2: (Color online) Controllable levels degeneracy in non-hermitian superconducting circuit. (a) Effective energy levels and (b) damping rates as the function of effective coupling $g_{e}$. The three colored curves correspond to the phase differences: $\Delta\theta=0.500\pi$ (blue curves), $0.501\pi$ (red curves), and $0.502\pi$ (black curves). The two qubits are assumed in the ground states $\langle\sigma^{j}_{z}\rangle=-1$, with $j=1,2$. (c) Effective energy levels as the function of resonator-qubit coupling amplitude $\lambda_{j}$. The four colored curves correspond to different parameters: $\Delta\theta=0.500\pi$, $\Gamma_{2}=\Gamma_{1}$, $g_{e}/2\pi=1.137$ MHz, and $\gamma_{a}/2\pi=65$ MHz (blue curves); $g_{e}/2\pi=1.224$ MHz (green curves); $\gamma_{a}/2\pi=66$ MHz (red curves); $\Gamma_{2}=0.99\Gamma_{1}$(black curves). $\langle\sigma^{(1)}_{z}\rangle=-1$, and the un-mentioned parameters in the green, red, and black curves are the same as in blue curves. (d) Effective energy levels as function of qubits’ populations $\langle\sigma^{j}_{z}\rangle$. The three colored curves correspond to different parameters: $\Delta\theta=0.500\pi$, $\lambda_{j}/2\pi=11$ MHz, $g_{e}/2\pi=1.137$ MHz (blue curves); $\Delta\theta=0.5005\pi$; $g_{e}/2\pi=1.224$ MHz, $\lambda_{j}/2\pi=11$ MHz. (black curves). $\langle\sigma^{j}_{z}\rangle=(-1)^{(j+1)}\langle\sigma_{z}\rangle$, the un-mentioned parameters in the red and black curves are the same as in blue curves. The other parameters in (a)-(d) are: $\omega_{a}/2\pi=4.475$ GHz, $\omega_{1}/2\pi=4.5$ GHz, $\omega_{2}/2\pi=4.505$ GHz, $\omega^{(max)}_{c}/2\pi=5.20$ GHz, $\gamma_{1}/2\pi=1.00$ MHz, $\gamma_{2}/2\pi=1.01$ MHz, $g_{xy}/2\pi=4.0$ MHz, $\lambda_{j}/2\pi=11$ MHz, $g_{1}/2\pi=30$ MHz, $g_{2}/2\pi=30.3$ MHz, and $\gamma_{a}/2\pi=65$ MHz.
In the rotating coordinates at resonator’s frequency $\omega_{a}$, the motion equations of cavity photon and qubit-$j$ can be written as $\dot{a}=-\gamma_{a}a-i\sum_{j=1,2}\lambda_{j}\exp{(-i\theta_{j})}\sigma^{(j)}_{-}$ and $\dot{\sigma}^{(j)}_{-}=-\left[\gamma_{j}/2+i\left(\Delta_{ja}+g^{2}_{j}/\Delta_{jc}\right)\right]\sigma^{(j)}_{-}-i\lambda_{j}\exp{(i\theta_{j})}\sigma^{(j)}_{z}a-ig_{e}\sigma^{(j)}_{z}\sigma^{(j+1)}_{-}$, with $\Delta_{ja}=\omega_{j}-\omega_{a}$ and $j+1=1$ for $j=2$. In the case of $\gamma_{a}\gg g_{j},\lambda_{j}\gg\gamma_{j}$ and $|\Delta_{ja}|/\gamma_{a}\approx 1/3$, with the adiabatic approximation, we can get $a=-i(\lambda_{1}/\gamma_{a})\sum_{j=1,2}\lambda_{j}\exp{(-i\theta_{j})}\sigma^{(j)}_{-}$. Substituting it into the motion equation of ladder operators, thus we get
@@
The Pauli-Z operator $\sigma^{(j)}_{z}$ describes the population of qubit-$j$ and its steady value can be tuned with external microwave signals. Though coupling to a high-loss resonator, the decoherence time of qubit should be much longer than the operation times (usually below 100 ns) for preparation and readout of single qubit’s quantum states especially for the tantalum-based superconducting qubit[42, 43]. Thus the population of qubit-$j$ can be considered as staying in the transient stability value $\langle\sigma^{(j)}_{z}\rangle$ during the measurement for the quantum state of single qubit, and its value can be tuned by the qubit-coupler coupling strength, and also the resonator-qubit coupling phases and amplitudes. Thus the effective Hamiltonian for non-Hermitian qubit-qubit coupling can be written as
@@
where $\Delta\theta=\theta_{1}-\theta_{2}$, $\Delta^{\prime}_{ja}=\Delta_{ja}+g^{2}_{j}/\Delta_{jc}$, and $\Gamma_{j}=\gamma_{j}-\lambda^{2}_{j}\langle\sigma^{(j)}_{z}\rangle/\gamma_{a}$, with $j=1,2$. The non-diagonal terms in Eq.(3) label total effective qubit-qubit coupling strength in the non-hermite superconducting circuit. The two Eigen-modes of $H^{\prime\prime}$ can be calculated as
@@
The real parts $\Re{(\omega_{\pm})}$ and imaginary parts $\Im{(\omega_{\pm})}$ describe the effective energy levels and effective damping rates of two qubits, respectively. The values of $R$ and $I$ affect the separations of complex energy spectrum between two Eigen-modes, and the signs of $R$ decide the level repulsions or attractions, while the sign of $I$ describes the relative size of two effective damping rates.
@@ III Controllable Non-Hermitian coupling
The experimental tunable parameters in this article include the resonator-qubit coupling phase and amplitude, effective qubit-qubit coupling, and occupation populations of qubits. If the parameters is tuned to satisfy $R<0$ and $I=0$, thus the $\sqrt{R+iI}$ becomes a pure imaginary number($\Delta E_{q}=0$), and the energy levels of two qubits merge as one and the exceptional point appears. Even for a tiny nonzero value for $I$, the energy levels of two qubits will misalign and the exceptional point should be absent[25].
@@
By setting two qubits in the ground states, the effective energy levels and damping rates of two Eigen-modes as the function effective qubit-qubit coupling ($g_{e}$) are shown in Figs. 2(a) and  2(b), respectively. The separations of two qubits’ effective energy levels (or damping rates) becomes narrower (or wider) in certain parameter regimes corresponding to the the energy level attraction. To see clearer, the curves in blue and black color shift up -0.5 MHz or 0.5 MHz in Fig. 2(a), respectively. In the case of $\Delta\theta=\pi/2$, the first term in the right side of Eq.(6) is always zero, since $\Delta^{\prime}_{1a}-\Delta^{\prime}_{2a}\neq 0$ in parameter area of drawing, so there is no level degenerate point on the blue color curves in Fig. 2(a) and the insert figure. For $\Delta\theta\neq\pi/2$, as the variations of $g_{c}$(or $\omega_{c}$), it is possible to satisfy $I=0$ so the level degenerate points appear on the pairs of two red or black color curves The scattering curves for damping rates of two qubits are shown in Fig. 2(b), to see clearer, the red and black color curves shift up 0.1 MHz and 0.2 MHz, respectively. The two branch of effective damping rates exchange with each other in two red or black color curves, and the corresponding exchange positions coincide well with those of level degenerate points in Fig. 2(a), which indicate the positions of exceptional points.
@@
Figure 3: (Color online) Controllable quantum state evolutions in energy level attraction regimes. (a) The energy level difference($\Delta E_{q}$) and (b) damping rate difference ($\Delta\gamma_{q}$) as functions of phase difference ($\Delta\theta$) and coupling strength ($g_{e}$). The two qubits are assumed as in the ground states, $\langle\sigma^{(j)}_{z}\rangle=-1$ with $j=1,2$. The evolutions of qubits’ excited state occupation probabilities for (c) qubit $1$ and (d) qubit $2$ The four colored curves correspond to different phase differences: $\Delta\theta=0$ (blue-solid curves); $\Delta\theta=\pi/3$ (green-dashed curves); $\Delta\theta=5\pi/12$ (red-dotted curves); $\Delta\theta=\pi/2$ (black-dash-dotted curves). The qubit $1$ and qubit $2$ respectively start in excited and ground states. A small effective coupling ($g_{e}/2\pi=-80$ kHz) is chosen for energy level attractions in (c) and (d). The other parameters of four images are: $\omega_{a}/2\pi=4.475$ GHz, $\omega_{1}/2\pi=4.5$ GHz, $\omega_{2}/2\pi=4.505$ GHz, $\omega^{(max)}_{c}/2\pi=5.20$ GHz, $\gamma_{1}/2\pi=1.00$ MHz, $\gamma_{2}/2\pi=1.01$ MHz, $g_{xy}/2\pi=4.0$ MHz, $\lambda_{j}/2\pi=11$ MHz, $g_{1}/2\pi=30$ MHz, $g_{2}/2\pi=30.3$ MHz, and $\gamma_{a}/2\pi=65$ MHz.
@@
The effects of resonator-qubit coupling amplitudes on energy levels of two qubit are studied in Fig. 2(c). In the case of $\Delta\theta=\pi/2$ and $\Gamma_{1}-\Gamma_{2}=0$, the first and second terms in Eq.(6) are always zeroes, so the energy level degeneracy appear for arbitrary value of $\lambda_{j}$ (energy level attraction regimes) in the blue curves. The change for the values of $g_{e}$ can shift the position of the Exception point and also the value of energy levels as indicated by the green curves (compared with the blue curves). If the $\Gamma_{1}-\Gamma_{2}\neq 0$, because of no solution for $\Delta^{\prime}_{1a}-\Delta^{\prime}_{2a}\neq 0$ in the parameter area of drawing, so the energy level degenerate points does not appear in the red and black curves. As indicated by the motion equation of $\sigma^{(j)}_{z}$, the effective energy levels of two qubits can be tuned by changing the quantum states of qubits.
@@
The free evolutions of quantum states under non-hermitian coupling can be calculated with the semiclassical limit method by disregarding the quantum jump terms[31, 47, 48]. By preparing the qubit 1 in the excited states and qubit 2 in the ground states, the time evolution of occupation probabilities for qubits’ excited state can be calculated with the semiclassical master equation $\dot{\rho}=-i(H^{\prime\prime}\rho-\rho H^{\prime\prime})$, where the $\rho$ is the density operator. To stay in the level attraction regimes, the effective qubit-qubit coupling $g_{e}$ is tuned to a small value, $g_{e}/2\pi=-80$ kHz. The time evolution of occupation probabilities for qubit-$1$ ($P^{(se)}_{1}$) starting in excited state and qubit-$2$ ($P^{(sg)}_{2}$) starting in ground state are shown in Fig. 3(c) and Fig. 3(d), respectively. The periodic oscillation of the curves originates from the nonzero frequency detuning $\Delta_{ja}$ in $H^{\prime\prime}$, and the effects of the resonator-qubit coupling phase can be reflected by the variation of the envelopes.